OpenStudy (anonymous):

Find a cubic function with the given zeros. -2, 5, -6

OpenStudy (anonymous):

start with the polynomial in factored form as \[(x+2)(x-5)(x+6)\]

OpenStudy (anonymous):

then if you have to, multiply out

OpenStudy (anonymous):

what do you mean ? like multiply together?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

first maybe \[(x-5)(x+6)=x^2+x-30\] and then \[(x+2)(x^2+x-30)=...\]

OpenStudy (anonymous):

oh, x^3+x-30x+2x^2+2x-60

OpenStudy (anonymous):

x^3+2x^2-27x-60 right?